1) Good X is elastic. Good Y is a substitute. You are considering increasing Px. Will your revenue rise or fall? Answer: the total revenue (joint TR for X and Y) will (pick one: rise, fall, remain the same, or ambiguous) because ...

2) Good X is elastic. Good Y is a substitute. You are considering decreasing Px. Will your revenue rise or fall? Answer: the total revenue (joint TR for X and Y) will (pick one: rise, fall, remain the same, or ambiguous) because ..

3) Good X is inelastic. Good Y is a substitute. You are considering increasing Px. Will your revenue rise or fall? Answer: the total revenue (joint TR for X and Y) will (pick one: rise, fall, remain the same, or ambiguous) because ...

4) Good X is inelastic. Good Y is a substitute. You are considering decreasing Px. Will your revenue rise or fall? Answer: the total revenue (joint TR for X and Y) will (pick one: rise, fall, remain the same, or ambiguous) because ...

5) Good X is elastic. Good Y is a complement. You are considering increasing Px. Will your revenue rise or fall? Answer: the total revenue (joint TR for X and Y) will (pick one: rise, fall, remain the same, or ambiguous) because ...

6) Good X is elastic. Good Y is a complement. You are considering decreasing Px. Will your revenue rise or fall? Answer: the total revenue (joint TR for X and Y) will (pick one: rise, fall, remain the same, or ambiguous) because ...

7) Good X is inelastic. Good Y is a complement. You are considering increasing Px. Will your revenue rise or fall? Answer: the total revenue for X and Y will (rise, fall, remain the same, or ambiguous) because ...

8) Good X is inelastic. Good Y is a complement. You are considering decreasing Px. Will your revenue rise or fall? Answer: the total revenue (joint TR for X and Y) will (pick one: rise, fall, remain the same, or ambiguous) because ...
Answer #1

(1) Ambiguous

Increase in Px will decrease the quantity demanded of X, and increase the demand for Y.

Since X is elastic, higher price of X will decrease TRx. Higher demand for Y will increase both price and quantity of Y. So TRy will increase. Therefore, net effect on total revenue (TR = TRx + TRy) cannot be precisely computed. TR will increase (decrease) if Decrease in TRx is less than (more than) the increase in TRy.

(2) Ambiguous

Decrease in Px will increase the quantity demanded of X, and decrease the demand for Y.

Since X is elastic, lower price of X will increase TRx. Lower demand for Y will decrease both price and quantity of Y. So TRy will decrease. Therefore, net effect on total revenue (TR = TRx + TRy) cannot be precisely computed. TR will increase (decrease) if Decrease in TRy is less than (more than) the increase in TRx.

(3) Increase

Increase in Px will decrease the quantity demanded of X, and increase the demand for Y.

Since X is inelastic, higher price of X will increase TRx. Higher demand for Y will increase both price and quantity of Y. So TRy will increase. Therefore, net effect on total revenue (TR = TRx + TRy) is an Increase.

(4) Decrease

Decrease in Px will increase the quantity demanded of X, and decrease the demand for Y.

Since X is inelastic, lower price of X will decrease TRx. Lower demand for Y will decrease both price and quantity of Y. So TRy will decrease. Therefore, net effect on total revenue (TR = TRx + TRy) is a Decrease.

Note: As HOMEWORKLIB Answering Policy, 1st 4 questions are answered.